# MMGRid: Navigating Temporal-aware and Cross-domain Generative Recommendation via Model Merging

### v1

@@ Abstract.
Model merging (MM) offers an efficient mechanism for integrating multiple specialized models without access to original training data or costly retraining. While MM has demonstrated success in natural language processing and computer vision, its role in recommender systems (RSs) remains largely unexplored. Recently, Generative Recommendation (GR) has emerged as a new paradigm in RSs, characterized by rapidly growing model scales and substantial computational costs, making MM particularly appealing for cost-sensitive deployment scenarios. In this work, we present the first systematic study of MM in GR through a contextual lens. We focus on a fundamental yet underexplored challenge in real-world GR: how to merge generative recommenders specialized to different real-world contexts, arising from temporal evolving user behaviors and heterogeneous application domains. To this end, we propose a unified framework MMGRid , a structured contextual grid of GR checkpoints that organizes models trained under diverse contexts induced by temporal evolution and domain diversity. All checkpoints are derived from a shared base large language model (LLM) but fine-tuned on context-specific data, forming a realistic and controlled model space for systematically analyzing MM across GR paradigms and merging algorithms. Our investigation reveals several key insights. First, training GR models from LLMs can introduce parameter conflicts during merging due to token distribution shifts and objective disparities; such conflicts can be alleviated by disentangling task-aware and context-specific parameter changes via base model replacement. Second, incremental training across contexts induces recency bias, which can be effectively balanced through weighted contextual merging. Notably, we observe that optimal merging weights correlate with context-dependent interaction characteristics, offering practical guidance for weight selection in real-world deployments. We open-source our framework, training pipelines, evaluation protocols, and contextual checkpoints to support future research on MM in GR.
@@ 1. Introduction
In machine learning, Model merging (MM) (Yang et al., 2024) has recently emerged as an efficient paradigm for integrating the capabilities of multiple specialized models into a single, universal model without requiring access to the original training data or additional expensive optimization. Building on the idea that fine-tuned models can be represented as parameter deltas from a shared pretrained backbone, MM directly operates on model parameters, typically through linear interpolation (Wortsman et al., 2022), task-vector arithmetic (Ilharco et al., 2022), or geometry-aware fusion (Yadav et al., 2023), to aggregate diverse features, skills, or domain knowledges. This property naturally aligns with modern machine learning pipelines, which follow the ubiquitous ”pre-training → post-training” paradigm (Kumar et al., 2025; Zhuang et al., 2021). Moreover, under the influence of scaling laws (Kaplan et al., 2020), increasing model sizes further exacerbate the cost of retraining from scratch. MM bypasses this costly process by directly merging post-trained models, offering a more efficient and scalable alternative to traditional training pipelines.
@@
The effectiveness of MM in reusing and recombining specialized expertise has led to its widespread adoption in computer vision (Ilharco et al., 2022; Wortsman et al., 2022), natural language processing (NLP) (Akiba et al., 2025), and multimodal learning (Aiello et al., 2023). Recommendation Systems (RSs), as an application-driven machine learning problem, are also encompassed within this broad paradigm. However, traditional RSs, characterized by multi-stage pipelines of heterogeneous components built on domain-specialized signals, have offered little incentive for parameter-level fusion (Deng et al., 2025b). This situation is fundamentally changing with the recent emergence of Generative Recommendation (GR) (Rajput et al., 2023; Tan et al., 2025; Chen et al., 2024; Zhai et al., 2024; Du et al., 2025). Inspired by the evolution of NLP towards Large Language Models (LLMs), GR replaces fragmented multi-stage architectures with end-to-end autoregressive modeling (Deng et al., 2025b), unifying representation learning, semantic reasoning, and decision making within a single generative framework. This architectural unification not only enhances semantic expressiveness and cross-scenario generalization, but also introduces a training and deployment pattern closely aligned with LLMs, making MM a technically viable yet previously unexplored strategy for GR. Furthermore, real-world RS deployments face persistent constraints, including stringent cost sensitivity (Zhou et al., 2025a) and the need for continual model adaptation driven by ever-evolving user behaviors and streaming interaction data (Guo et al., 2025; Yoo et al., 2025). MM, with its ability to efficiently recombine specialized checkpoints, is uniquely positioned to address such contextual variation. However, despite this strong practical motivation, a systematic understanding of how MM behaves under realistic contextual shifts in GR remains missing.
@@
In this work, we present the first systematic investigation of MM for GR from a contextual perspective. We introduce a unified experimental framework, MMGRid , which constructs a contextual grid as a structured model space composed of GR checkpoints specialized to different real-world contexts. The framework encompasses three representative GR paradigms: text-grounded generation (Bao et al., 2025), semantic ID modeling (Deng et al., 2025b; Zheng et al., 2024), and semantic embedding modeling (Chen et al., 2024), and two mainstream merging algorithms: weight-based merging (Ilharco et al., 2022) and subspace-based merging (Yadav et al., 2023), together with consistent training and inference pipelines. Building upon this contextual model space, we study MM along two fundamental dimensions of real-world context: heterogeneous recommendation domains and temporal evolution of user interactions. Rather than treating these as isolated settings, we view them as complementary sources of contextual variation that jointly shape the deployment landscape of GR. This perspective enables a unified analysis of merging behaviors across contexts that naturally arise in cross-domain recommendation and continual system updating.
@@
These explorations provide important insights into the applicability and challenges of MM for GR, revealing both the opportunities and limitations of parameter-level merging in recommendation scenarios. We believe that our findings will inform the design of more effective merging strategies tailored to the unique characteristics of GR models. Additionally, to promote reproducibility and encourage further investigation, we open-source our experimental framework  [^1] https://github.com/Joinn99/MMGRid, trained models, and evaluation protocols, which we hope will serve as a foundation for future research in this emerging direction.
@@ 2.1. Generative Recommendation
2.2. Model Merging
Model Merging (MM) is an effective technique that merges the parameters of multiple distinct models to create a single unified model with various objectives. A significant advantage of MM is its efficiency: it bypasses the substantial computational expense and the need for original training data associated with retraining from scratch (Wortsman et al., 2022; Ilharco et al., 2022). The merging is performed directly at the parameter level, resulting in a single model for inference. Here, we define the parameter-wise model merging problem as follows:
@@ 2.2. Model Merging
Assume we have $N$ models, $\Phi_{\Theta^{(1)}},\dots,\Phi_{\Theta^{(N)}}$, which share the same architecture. These models are typically fine-tuned from a common pre-trained model $\Phi_{\Theta}$. The parameters (or weights) of the $n$-th model are denoted as $\Theta^{(n)}$. The objective of model merging is to find a function $\textsc{merge}(\cdot)$ that combines the parameters of these models to produce a new set of parameters $\Theta^{*}$:
@@
Given the general formulation above, a variety of concrete task-vector MM methods for the function $\textsc{merge}(\cdot)$ have been proposed, which can be broadly classified into weight-based and subspace-based approaches.
@@ 3. Framework Design
In GR, a significant characteristic is the manifestation of scaling laws (Deng et al., 2025b). The improved generalization capacity of larger-scale models allows them to effectively integrate knowledge from various domains. However, this increase in model size concurrently elevates the cost of model tuning for domain merging and incremental updates, posing a significant challenge to cost-sensitive recommendation scenarios with real-world contexts like diverse application domains and temporal evolving of interactions. To address these challenges, MM techniques offer a promising solution by efficiently integrating the capabilities of multiple models through parameter space fusion without requiring additional training. In this work, we seek to create a experimantal framework, named MMGRid , to explore the potential of MM in GR, identifying hidden issues and challenges and promising directions for solutions.
@@ 3.1. Identifying Real-world Contexts of GR
Suppose we have models $\Phi_{\Theta^{(\mathcal{D}_{1},t_{1})}},\dots,\Phi_{\Theta^{(\mathcal{D}_{N},t_{1})}}$, which are trained on different $N$ domains $\mathcal{D}_{1},\dots,\mathcal{D}_{N}$ at time $t_{1}$ respectively. The objective of cross-domain merging is to find a function $\textsc{Merge}(\cdot)$ that combines the parameters of these models to produce a new set of parameters $\Theta^{*}$:
@@ 3.2. Creating Contextual Checkpoint Grid
In the context of GR, several factors introduce key distinctions from typical application domains of MM. The temporal nature of user interaction data, parametric variations across different types of GR approaches, and differences in backbone model architectures collectively pose new challenges for model merging in RSs. In this work, we aim to closely approximate realistic RS scenarios and their practical requirements for model merging. Therefore, we construct a unified experimental framework that enables the creation of different types of GR models across various domains and time periods, all built upon a unified backbone, facilitating MM experiments under conditions that closely mirror real-world scenarios and align with practical demands.
@@ 4. Creating Unified GR via Cross-domain Merging: Challenges, Causes and Solutions
In this section, following the emerging trend of developing unified GR models, we elaborate on the challenges faced when creating a cross-domain GR through MM. We systematically discuss key difficulties and potential sources that affect the performance of cross-domain merging, and provide potential practical solutions to address these issues.
@@ 4.1. Initial Experiments and Analysis
To explore the potential of MM on cross-domain GR, we first adopt the most fundamental setting: merging between two domains and evaluating performance of the merged model on each source domain separately.
@@
Finding 1: Text-Grounded GR shows semantic correlated positive transfer. In the initial experiments, Text-grounded is the only GR paradigm that achieves performance improvements through MM, and the magnitude of improvement appears correlated with the semantic characteristics of domains. Domains with diverse semantic knowledge, such as Books and Movies, tend to provide greater benefits when merged with other domains. Domain pairs with totally different semantic characteristics, such as Cellphones and Sports, tend to show performance degradation. This suggests that text-grounded representations can effectively leverage cross-domain semantic commonalities encoded in the LLM’s pretrained knowledge.
@@ 4.3. Disentangling Task-Aware Parameter Shift
Task vectors for models at $t_{2}$ are then recomputed based on $\Theta^{(\mathcal{D}_{i}\cup\mathcal{D}_{j},t)}$ before merging. Figure 4 presents the merging performance. Results demonstrate further performance improvements with the neutral base strategy, suggesting that this approach more effectively separates domain-specific knowledge, thereby reducing redundant task-aware information in task vectors. It is worth noting that similar strategies have been discussed in the literature of MM (Choi et al., 2025), which corroborates our conclusions.
@@ 5. Balancing Evolving User Preferences via Temporal Merging
In this section, we are seeking to exploit the potential of MM in incremental training scenraio of practical RSs. In traditional applications of MM, merging checkpoints from the training process to facilitate fine-tuning is a widely-adopted practice (Liu et al., 2024). While this shares similarities with continual training in RSs, a critical distinction exists: continual training data in RSs inherently encodes the temporal variance of distribution in the training data. In incrementally trained RSs, newly incorporated training data originates from fresh user interactions, typically encoding recent preference shifts among currently active users (Guo et al., 2025). This suggests that parameter changes between old and new checkpoints may implicitly capture trends in evolving user preferences, hinting at the possibility of leveraging these trends to fine-tune RSs and balance historical and emerging preferences to achieve better overall performance.
@@
In this section, we investigate how to perform such optimization through MM of temporal checkpoints. Specifically, we examine whether merging models from different time periods can yield a unified model that outperforms either individual checkpoint by effectively balancing temporal dynamics in user preferences.
@@ 5.2. Balancing Evolving Preferences via Weighted-based Merging
The above preliminary analysis demonstrates the potential of MM in balancing historical and emerging user preferences in the context of incremental training for RSs. By leveraging cross-domain statistics, we can effectively guide the merging process to achieve better performance than naive approaches. However, several limitations remain. First, the linear regression model, while simple and interpretable, may not capture complex nonlinear relationships between domain characteristics and optimal merging weights. Second, our analysis focuses primarily on item recency as a domain characteristic, but other factors such as user engagement patterns, item diversity, and seasonal trends may also play important roles.

### v2

@@ Abstract.
Model merging ($M$) offers an efficient mechanism for integrating multiple specialized models without access to original training data or costly retraining. While $M$ has demonstrated success in natural language processing and computer vision, its role in recommender systems (RSs) remains largely unexplored. Recently, Generative Recommendation (GR) has emerged as a new paradigm in RSs, characterized by rapidly growing model scales and substantial computational costs, making $M$ particularly appealing for cost-sensitive deployment scenarios. In this work, we present the first systematic study of $M$ in GR through a contextual lens. We focus on a fundamental yet underexplored challenge in real-world GR: how to merge generative recommenders specialized to different real-world contexts, arising from temporal evolving user behaviors and heterogeneous application domains. To this end, we propose a unified framework MMGRid , a structured contextual grid of GR checkpoints that organizes models trained under diverse contexts induced by temporal evolution and domain diversity. All checkpoints are derived from a shared base large language model (LLM) but fine-tuned on context-specific data, forming a realistic and controlled model space for systematically analyzing $M$ across GR paradigms and merging algorithms. Our investigation reveals several key insights. First, training GR models from LLMs can introduce parameter conflicts during merging due to token distribution shifts and objective disparities; such conflicts can be alleviated by disentangling task-aware and context-specific parameter changes via base model replacement. Second, incremental training across contexts induces recency bias, which can be effectively balanced through weighted contextual merging. Notably, we observe that optimal merging weights correlate with context-dependent interaction characteristics, offering practical guidance for weight selection in real-world deployments. We open-source our framework, training pipelines, evaluation protocols, and contextual checkpoints to support future research on $M$ in GR.
@@ 1. Introduction
In machine learning, Model merging ($M$) (Yang et al., 2024) has recently emerged as an efficient paradigm for integrating the capabilities of multiple specialized models into a single, universal model without requiring access to the original training data or additional expensive optimization. Building on the idea that fine-tuned models can be represented as parameter deltas from a shared pretrained backbone, $M$ directly operates on model parameters, typically through linear interpolation (Wortsman et al., 2022), task-vector arithmetic (Ilharco et al., 2022), or geometry-aware fusion (Yadav et al., 2023), to aggregate diverse features, skills, or domain knowledges. This property naturally aligns with modern machine learning pipelines, which follow the ubiquitous ”pre-training → post-training” paradigm (Kumar et al., 2025; Zhuang et al., 2021). Moreover, under the influence of scaling laws (Kaplan et al., 2020), increasing model sizes further exacerbate the cost of retraining from scratch. $M$ bypasses this costly process by directly merging post-trained models, offering a more efficient and scalable alternative to traditional training pipelines.
@@
The effectiveness of $M$ in reusing and recombining specialized expertise has led to its widespread adoption in computer vision (Ilharco et al., 2022; Wortsman et al., 2022), natural language processing (NLP) (Akiba et al., 2025), and multimodal learning (Aiello et al., 2023). Recommendation Systems (RSs), as an application-driven machine learning problem, are also encompassed within this broad paradigm. However, traditional RSs, characterized by multi-stage pipelines of heterogeneous components built on domain-specialized signals, have offered little incentive for parameter-level fusion (Deng et al., 2025b). This situation is fundamentally changing with the recent emergence of Generative Recommendation (GR) (Rajput et al., 2023; Tan et al., 2025; Chen et al., 2024; Zhai et al., 2024; Du et al., 2025). Inspired by the evolution of NLP towards Large Language Models (LLMs), GR replaces fragmented multi-stage architectures with end-to-end autoregressive modeling (Deng et al., 2025b), unifying representation learning, semantic reasoning, and decision making within a single generative framework. This architectural unification not only enhances semantic expressiveness and cross-scenario generalization, but also introduces a training and deployment pattern closely aligned with LLMs, making $M$ a technically viable yet previously unexplored strategy for GR. Furthermore, real-world RS deployments face persistent constraints, including stringent cost sensitivity (Zhou et al., 2025a) and the need for continual model adaptation driven by ever-evolving user behaviors and streaming interaction data (Guo et al., 2025; Yoo et al., 2025). $M$, with its ability to efficiently recombine specialized checkpoints, is uniquely positioned to address such contextual variation. However, despite this strong practical motivation, a systematic understanding of how $M$ behaves under realistic contextual shifts in GR remains missing.
@@
In this work, we present the first systematic investigation of $M$ for GR from a contextual perspective. We introduce a unified experimental framework, MMGRid , which constructs a contextual grid as a structured model space composed of GR checkpoints specialized to different real-world contexts. The framework encompasses three representative GR paradigms: text-grounded generation (Bao et al., 2025), semantic ID modeling (Deng et al., 2025b; Zheng et al., 2024), and semantic embedding modeling (Chen et al., 2024), and two mainstream merging algorithms: weight-based merging (Ilharco et al., 2022) and subspace-based merging (Yadav et al., 2023), together with consistent training and inference pipelines. Building upon this contextual model space, we study $M$ along two fundamental dimensions of real-world context: heterogeneous recommendation domains and temporal evolution of user interactions. Rather than treating these as isolated settings, we view them as complementary sources of contextual variation that jointly shape the deployment landscape of GR. This perspective enables a unified analysis of merging behaviors across contexts that naturally arise in cross-domain recommendation and continual system updating.
@@
These explorations provide important insights into the applicability and challenges of $M$ for GR, revealing both the opportunities and limitations of parameter-level merging in recommendation scenarios. We believe that our findings will inform the design of more effective merging strategies tailored to the unique characteristics of GR models. Additionally, to promote reproducibility and encourage further investigation, we open-source our experimental framework  [^1] https://github.com/Joinn99/MMGRid, trained models, and evaluation protocols, which we hope will serve as a foundation for future research in this emerging direction.
@@ 2.1. Generative Recommendation
$2.$ Model Merging
Model Merging ($M$) is an effective technique that merges the parameters of multiple distinct models to create a single unified model with various objectives. A significant advantage of $M$ is its efficiency: it bypasses the substantial computational expense and the need for original training data associated with retraining from scratch (Wortsman et al., 2022; Ilharco et al., 2022). The merging is performed directly at the parameter level, resulting in a single model for inference. Here, we define the parameter-wise model merging problem as follows:
@@
Assume we have NN models, $\Phi_{\Theta^{(1)}},\dots,\Phi_{\Theta^{(N)}}$, which share the same architecture. These models are typically fine-tuned from a common pre-trained model $\Phi_{\Theta}$. The parameters (or weights) of the $n$-th model are denoted as $\Theta^{(n)}$. The objective of model merging is to find a function $\textsc{merge}(\cdot)$ that combines the parameters of these models to produce a new set of parameters $\Theta^{*}$:
@@
Given the general formulation above, a variety of concrete task-vector $M$ methods for the function $\textsc{merge}(\cdot)$ have been proposed, which can be broadly classified into weight-based and subspace-based approaches.
@@ 3. Framework Design
In GR, a significant characteristic is the manifestation of scaling laws (Deng et al., 2025b). The improved generalization capacity of larger-scale models allows them to effectively integrate knowledge from various domains. However, this increase in model size concurrently elevates the cost of model tuning for domain merging and incremental updates, posing a significant challenge to cost-sensitive recommendation scenarios with real-world contexts like diverse application domains and temporal evolving of interactions. To address these challenges, $M$ techniques offer a promising solution by efficiently integrating the capabilities of multiple models through parameter space fusion without requiring additional training. In this work, we seek to create a experimantal framework, named MMGRid , to explore the potential of $M$ in GR, identifying hidden issues and challenges and promising directions for solutions.
@@ 3.1. Identifying Real-world Contexts of GR
Suppose we have models $\Phi_{\Theta^{(\mathcal{D}_{1},t_{1})}},\dots,\Phi_{\Theta^{(\mathcal{D}_{N},t_{1})}}$, which are trained on different NN domains $\mathcal{D}_{1},\dots,\mathcal{D}_{N}$ at time $t_{1}$ respectively. The objective of cross-domain merging is to find a function $\textsc{Merge}(\cdot)$ that combines the parameters of these models to produce a new set of parameters $\Theta^{*}$:
@@ 3.2. Creating Contextual Checkpoint Grid
In the context of GR, several factors introduce key distinctions from typical application domains of $M$. The temporal nature of user interaction data, parametric variations across different types of GR approaches, and differences in backbone model architectures collectively pose new challenges for model merging in RSs. In this work, we aim to closely approximate realistic RS scenarios and their practical requirements for model merging. Therefore, we construct a unified experimental framework that enables the creation of different types of GR models across various domains and time periods, all built upon a unified backbone, facilitating $M$ experiments under conditions that closely mirror real-world scenarios and align with practical demands.
@@ 4. Creating Unified GR via Cross-domain Merging: Challenges, Causes and Solutions
In this section, following the emerging trend of developing unified GR models, we elaborate on the challenges faced when creating a cross-domain GR through $M$. We systematically discuss key difficulties and potential sources that affect the performance of cross-domain merging, and provide potential practical solutions to address these issues.
@@ 4.1. Initial Experiments and Analysis
To explore the potential of $M$ on cross-domain GR, we first adopt the most fundamental setting: merging between two domains and evaluating performance of the merged model on each source domain separately.
@@
Finding 1: Text-Grounded GR shows semantic correlated positive transfer. In the initial experiments, Text-grounded is the only GR paradigm that achieves performance improvements through $M$, and the magnitude of improvement appears correlated with the semantic characteristics of domains. Domains with diverse semantic knowledge, such as Books and Movies, tend to provide greater benefits when merged with other domains. Domain pairs with totally different semantic characteristics, such as Cellphones and Sports, tend to show performance degradation. This suggests that text-grounded representations can effectively leverage cross-domain semantic commonalities encoded in the LLM’s pretrained knowledge.
@@ 4.3. Disentangling Task-Aware Parameter Shift
Task vectors for models at $t_{2}$ are then recomputed based on $\Theta^{(\mathcal{D}_{i}\cup\mathcal{D}_{j},t)}$ before merging. Figure 4 presents the merging performance. Results demonstrate further performance improvements with the neutral base strategy, suggesting that this approach more effectively separates domain-specific knowledge, thereby reducing redundant task-aware information in task vectors. It is worth noting that similar strategies have been discussed in the literature of $M$ (Choi et al., 2025), which corroborates our conclusions.
@@ 5. Balancing Evolving User Preferences via Temporal Merging
In this section, we are seeking to exploit the potential of $M$ in incremental training scenraio of practical RSs. In traditional applications of $M$, merging checkpoints from the training process to facilitate fine-tuning is a widely-adopted practice (Liu et al., 2024). While this shares similarities with continual training in RSs, a critical distinction exists: continual training data in RSs inherently encodes the temporal variance of distribution in the training data. In incrementally trained RSs, newly incorporated training data originates from fresh user interactions, typically encoding recent preference shifts among currently active users (Guo et al., 2025). This suggests that parameter changes between old and new checkpoints may implicitly capture trends in evolving user preferences, hinting at the possibility of leveraging these trends to fine-tune RSs and balance historical and emerging preferences to achieve better overall performance.
@@
In this section, we investigate how to perform such optimization through $M$ of temporal checkpoints. Specifically, we examine whether merging models from different time periods can yield a unified model that outperforms either individual checkpoint by effectively balancing temporal dynamics in user preferences.
@@ 5.2. Balancing Evolving Preferences via Weighted-based Merging
The above preliminary analysis demonstrates the potential of $M$ in balancing historical and emerging user preferences in the context of incremental training for RSs. By leveraging cross-domain statistics, we can effectively guide the merging process to achieve better performance than naive approaches. However, several limitations remain. First, the linear regression model, while simple and interpretable, may not capture complex nonlinear relationships between domain characteristics and optimal merging weights. Second, our analysis focuses primarily on item recency as a domain characteristic, but other factors such as user engagement patterns, item diversity, and seasonal trends may also play important roles.